OpenStudy (anonymous):

What is the least common multiple of 6 and 3?

OpenStudy (anonymous):

hi again lol ill give u best answer if u get it first

OpenStudy (anonymous):

12

OpenStudy (anonymous):

is wrong

Parth (parthkohli):

\[\begin{array}{l|r}3&6,3\\\hline2&2,1\\\hline &1,1 \end{array} \\ \\ LCM(6,3) = 3 \times 2 \]

OpenStudy (anonymous):

12 is wrong its not on ma answrers

OpenStudy (anonymous):

Least common multiple of 3 and 6 is 6 haha xD

OpenStudy (tyteen4a03):

LCM of 6 and 3 is 6, because 3 * 2 = 6.

OpenStudy (anonymous):

thx guys

Parth (parthkohli):

lol, wow.\[6 \times 1 = 6\\3\times 2 = 6 \]Yes, it's 6.

OpenStudy (anonymous):

hhmm it can be also 6.

OpenStudy (anonymous):

but it cant b 12

OpenStudy (anonymous):

hhmmm!!!

OpenStudy (anonymous):

lol